OpenStudy (anonymous):

Use the definition of the derivative to prove the formula (f/g)' = (f'g-g'f)/g^2

OpenStudy (hba):

its the u/v form

OpenStudy (anonymous):

yes

OpenStudy (hba):

n if u want to check its right or wrong use the f.p.m

OpenStudy (jamesj):

So by definition of derivative, \[ (f/g)'(x) = \lim_{x \rightarrow a} \frac{ (f/g)(x) - (f/g)(a)}{x-a} \] Are you using this definition, or the one where the limit of delta x or h --> 0?

OpenStudy (hba):

no not abinatios method ur write james

OpenStudy (jamesj):

Lymeng_Chhorn, talk to me. Is what I've written down the definition of derivative you are using or not? Once we have that, I can show you the next steps. But be quick in responding as there are other people's questions out there as well.

OpenStudy (anonymous):

the limit of delta --> 0

OpenStudy (jamesj):

Ok, so consider the difference quotient \[ \frac{(f/g)(x + \Delta x) - (f/g)(x)}{\Delta x} \] This is equal to \[ \frac{f(x + \Delta x)/g(x+\Delta x)- f(x)/g(x)}{\Delta x} \] by definition of the quotient of functions. Now we know what we want this to be equal to. So you want to manipulate this expression somehow so you have a limit such as....

OpenStudy (jamesj):

\[ \frac{g(x) ( f(x+\Delta x) - f(x) )}{g(x)^2\Delta x} - f(x)\frac{g(x+\Delta x) - g(x)}{g(x)^2 \Delta x}\]

OpenStudy (jamesj):

Hence manipulate the difference quotient until you get it in a form like this and then take the limit.

OpenStudy (anonymous):

ok i got it :) Thank @JamesJ